# Relativistic Quantum Simulation of Hydrogen Sulfide for Hydrogen Energy via Hybrid Quantum-Classical Algorithms

### v1

@@ 2.1 Relativistic Electronic Structure Theory
Accurate modeling of heavy-element molecules requires the inclusion of relativistic effects in electronic structure calculations. For systems involving atoms such as sulfur, relativistic contributions—particularly scalar relativistic corrections and spin–orbit coupling—can significantly impact orbital energies and molecular properties.
The Dirac–Coulomb Hamiltonian provides a first-principles relativistic formulation of the molecular electronic structure. For a single electron in the field of a point nucleus, the Dirac equation is given by:
where is the speed of light, and are the Dirac matrices, is the momentum operator, and is the nuclear Coulomb potential. For a many-electron system, the Dirac–Coulomb Hamiltonian reads:
This four-component formalism introduces spinor wavefunctions and accounts for relativistic effects in a non-perturbative manner.
In practice, relativistic integrals are computed using quantum chemistry software such as DIRAC [12], which performs molecular mean-field calculations under the Dirac–Hartree–Fock approximation. These one- and two-electron integrals are then transformed into a second-quantized fermionic Hamiltonian.
@@ 4.3 Hydrogen Sulfide (H2S): Relativistic vs Non-Relativistic
Furthermore, we analyzed the number of measured Pauli strings and circuit depth under both encoding schemes. The relativistic Hamiltonians generally result in more Pauli terms, leading to increased measurement overhead (up to 15% more).
@@ Acknowledgements
This research was supported by the National Science and Technology Council (NSTC), R.O.C., under the research project ”Investigating Quantum Channels in Noisy Quantum Systems: Some Mathematical Problems in Quantum Information Theory” (Project No. 113-2115-M-A49-015-MY2). The authors would also like to acknowledge the use of AI-based language assistance tools for refining the English presentation of this manuscript.

### v2

@@ 2 Theoretical Framework
This section outlines the theoretical foundations upon which our study is based, particularly focusing on applying relativistic quantum chemistry to the electronic structure of the hydrogen sulfide (H2S) molecule. We begin by introducing the relativistic treatment of molecular systems, followed by describing how these concepts are encoded and simulated on hybrid quantum-classical computing architectures.
@@ 2.1 Relativistic Electronic Structure Theory
Accurate modeling of molecular systems containing heavy elements requires the explicit inclusion of relativistic effects in electronic structure calculations. In particular, for molecules involving atoms such as sulfur, relativistic contributions—especially scalar relativistic corrections and spin–orbit coupling—can significantly influence orbital energies, electron densities, and molecular properties.
A first-principles relativistic description of the electronic structure is provided by the Dirac–Coulomb Hamiltonian. For a single electron in the field of a point nucleus, the Dirac equation is given by
where denotes the speed of light, and are the Dirac matrices, is the momentum operator, and represents the nuclear Coulomb potential. Note that represents a three-component vector of matrices , and is written in bold to indicate its vectorial character. In contrast, is a single Dirac matrix associated with the mass term and is not boldfaced.
For an -electron system, the Dirac–Coulomb Hamiltonian takes the form
where the second term accounts for electron–electron Coulomb repulsion. This four-component formalism introduces spinor wavefunctions and incorporates relativistic effects in a nonperturbative manner.
In practice, relativistic molecular integrals are evaluated using quantum chemistry software packages such as DIRAC [12], which perform mean-field calculations based on the Dirac–Hartree–Fock (DHF) approximation. The resulting one-electron and two-electron integrals, expressed in a relativistic basis, are subsequently transformed into a second-quantized fermionic Hamiltonian suitable for quantum simulation.
@@ 3 Computational Methodology
This section outlines the computational strategy employed to simulate the electronic structure of hydrogen sulfide using a hybrid quantum-classical framework. Building upon the relativistic formulation established in the previous section, we translate the Dirac–Coulomb-based molecular Hamiltonian into a representation amenable to quantum simulation. This involves a sequence of classical preprocessing steps, quantum circuit design, and iterative variational procedures that collectively enable ground-state energy estimation on near-term quantum devices.
@@ 4 Results and Discussion
This section presents the simulation results for the molecular systems studied using the relativistic hybrid quantum-classical framework described in the previous sections. We begin with the hydrogen molecule (H2) as a benchmark system to validate the overall workflow and compare relativistic and non-relativistic integrals’ outputs. We then proceed to the primary target of this study, hydrogen sulfide (H2S), whose electronic structure and dissociation profile are analyzed in detail. We evaluate the accuracy of variational quantum eigensolver (VQE) results for each system against classical reference methods and visualize the corresponding potential energy surfaces.
@@ 4.3 Hydrogen Sulfide (H2S): Relativistic vs Non-Relativistic
Furthermore, we analyzed the number of measured Pauli strings and circuit depth under both encoding schemes. The relativistic Hamiltonians generally result in more Pauli terms, leading to increased measurement overhead (up to more).
@@ Acknowledgements
This research was supported by the National Science and Technology Council (NSTC), Taiwan, under the research project ”Investigating Quantum Channels in Noisy Quantum Systems: Some Mathematical Problems in Quantum Information Theory” (Project No. 113-2115-M-A49-015-MY2). The authors would also like to acknowledge the use of AI-based language assistance tools for refining the English presentation of this manuscript.